OpenStudy (atlchic):

Put 2x+5x=7y+2 in standard form.

OpenStudy (lgbasallote):

i believe the standard form is \[\large ax + by + c = 0\]

OpenStudy (atlchic):

yea, can u show me how 2 put that in standard form

OpenStudy (lgbasallote):

combine all the x first

OpenStudy (atlchic):

how show me on the screen plz

OpenStudy (lgbasallote):

2x + 5x <---what is 2 + 5?

OpenStudy (atlchic):

7

OpenStudy (lgbasallote):

so 2x + 5x is...?

OpenStudy (atlchic):

9y

OpenStudy (lgbasallote):

9y?

OpenStudy (lgbasallote):

2x + 5x is just like 2 + 5...just with an x

OpenStudy (atlchic):

7, when u add the y side u get 9

OpenStudy (lgbasallote):

you cant add x and y

OpenStudy (atlchic):

7 is 2+5

OpenStudy (anonymous):

2x+5x=7 7x+7y+2=0

OpenStudy (anonymous):

-7y=-7x+2 i think

OpenStudy (anonymous):

i think

OpenStudy (lgbasallote):

so if 2 + 5 is 7...what is 2x + 5x?

OpenStudy (atlchic):

7x

OpenStudy (lgbasallote):

good so 7x = 7y +2 only thing left now is to change in ax + by + c = 0 to do that...subtract both sides by 7y to bring it to the left side

OpenStudy (atlchic):

how do you do that show me plz

OpenStudy (lgbasallote):

7x - 7y = 7y + 2 - 7y

OpenStudy (lgbasallote):

note that our objective is to turn the right side into 0

OpenStudy (atlchic):

yea and

OpenStudy (lgbasallote):

so what's 7y - 7y?

OpenStudy (atlchic):

0

OpenStudy (lgbasallote):

good... 7x - 7y = 2 now what do you think comes next?

OpenStudy (atlchic):

subtract

OpenStudy (lgbasallote):

right..what do we subtract?

OpenStudy (atlchic):

2

OpenStudy (lgbasallote):

good...so the standard form is 7x - 7y -2 = 0

OpenStudy (anonymous):

yeah

OpenStudy (atlchic):

o yay some more questions

OpenStudy (atlchic):

is (-2,4), (1,3), (5,2), (1,4) A FUNCTION?